# Dynamical regimes of thermally convective emulsions

### v1

@@ Abstract
Emulsions are paramount in various interdisciplinary topical areas, yet a satisfactory understanding of their behavior in buoyancy-driven thermal flows has not been established. In the present work, we unravel the dynamical regimes of thermal convection in emulsions by leveraging a large set of mesoscale numerical simulations. Emulsions are prepared with a given volume fraction of the initially dispersed phase, $\phi$, ranging from dilute (low values of $\phi$) to jammed emulsions (high values of $\phi$), resulting in different rheological responses, i.e., from Newtonian to non-Newtonian yield-stress behaviors, respectively. We then characterize the dynamics of the emulsions in the paradigmatic setup of Rayleigh-Bénard convection, i.e., when confined between two parallel walls at different temperatures under the effect of buoyancy forces, the latter encoded in the dimensionless Rayleigh number $\operatorname{Ra}$. We thoroughly investigated the dynamics of the emulsion in the changing of $\phi$ and $\operatorname{Ra}$. For a given $\phi$, at increasing $\operatorname{Ra}$, we observe that the emulsion exhibits convection states, where structural changes may appear (i.e., droplet breakup, coalescence, or phase-inversion), which inevitably impact the emulsion rheology. For sufficiently high values of $\operatorname{Ra}$, two states of convection are observed: for low/moderate values of $\phi$ (Newtonian emulsions), we observe breakup-dominated dynamics, whereas, for high values of $\phi$ (non-Newtonian emulsions), we observe phase-inverted states. For both scenarios, the droplet size distribution depends on $\operatorname{Ra}$, and scaling laws for the average droplet size are analyzed and quantified. Our results offer unprecedented insights into the rich dynamics of emulsions under thermal convection, offering the first detailed characterization of the various dynamical regimes to be expected and their relation with structural changes occurring in such complex fluids.
@@ I Introduction
Emulsions are heterogeneous systems formed by a dispersion of droplets of a liquid phase (e.g., oil) in another immiscible liquid (e.g., water). Adding an emulsifier (e.g., surfactants) stabilizes the dispersion by hindering the coalescence of droplets [1, 2, 3, 4]. Owing to the presence of interfaces that can store part of the injected energy via elastic deformations [5, 6, 7, 8, 9], an emulsion exhibits a variety of rheological responses when stimulated with an external force. An important parameter – discriminating different rheological responses of the emulsion – is the volume fraction, $\phi$, i.e., the ratio between the volume of the dispersed phase and the total volume, which is related to the total interface area for a given number of droplets. For low values of $\phi$, the system behaves as a Newtonian fluid presenting an augmented effective viscosity that depends on $\phi$ and the viscosity ratio of the two phases [10, 11, 12]. However, for higher values of $\phi$, the rheology of the system is non-Newtonian in that it exhibits a shear-thinning rheology [13], which changes to yield stress behavior when $\phi$ is high enough [1, 14, 6, 7]. If the system is solicited with a sufficiently large external force, droplets can break up and coalesce [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], thus inducing fluctuations of the total interface area and, hence, modifying the rheological response of the system. The application of flowing emulsions in different manufacturing contexts, such as pharmaceutical [28], food [29] or energy industries [30], made these materials the subject of intense scrutiny for decades since a precise control of their stability and rheological properties under specific external forcing and boundary conditions is required in such contexts [31, 32]. Flowing emulsions exhibit highly complex features. Hence, numerical simulations are desirable tools of investigation to address regimes and processes particularly difficult to reproduce in experiments, thus unraveling multiple relevant questions. This aspect is witnessed by the extensive literature on computational studies on droplet/emulsion dynamics, with various methods developed [33, 34, 30, 35, 36, 37]. However, if, from one side, considerable effort has been made to investigate the behavior of emulsions under shear, pressure-driven, or even more complex flows (e.g., turbulence), from the other side, a satisfactory understanding of the behavior of emulsions in buoyancy-driven thermal flows remains incomplete. Thermally driven emulsions bear relevance in geophysics, for instance in the earth’s mantle convection [38, 39, 40, 41, 42], where convection currents transport heat from the interior to the planet’s surface, and explosive and nonexplosive lava flows [43, 44] or for analog modeling of geodynamic processes [45, 46, 47]. Also, petrochemical [48] and food cooking [49] applications are worth to be mentioned.
@@
Some computational studies considered droplets (or also bubbles) in convective flows, but they pertain to the case of convective multiphase flows with point-like dispersed objects [50, 51], or multiphase flows where finite-sized droplets/bubbles undergo a dynamics without any stabilization mechanism against their coalescence [52, 53, 54, 24, 55, 56, 57, 58], thus resulting in a system with overall Newtonian rheology. On the other hand, some theoretical/numerical works investigated the role played by the non-Newtonian rheology – and in particular the yield-stress rheology – on the onset of thermal convection using effective single fluid models  [59, 60, 61, 62, 63, 64, 65, 14, 66, 67]. These models rely on the use of the equations of continuum mechanics, supplemented with some non-trivial constitutive relations between the stress and the shear rate to account for non-Newtonian rheology. From the analysis of these effective single fluid models with yield-stress rheology, it has been found that conductive states become linearly stable in the presence of a finite yield stress [59, 60, 14]. Moreover, the system can enter a convective state only after applying a perturbation above a given intensity threshold. The latter increases upon approaching the critical Rayleigh number, marking the transition from conduction to convection [59]. Effective single fluid models, however, neglect the size of microscopic constituents (droplets/bubbles); hence, they cannot account for plasticity and structural changes at mesoscales [68, 69]; the latter, in turn, may have a non-trivial correlation with thermal plumes, as evidenced by some experiments on yield-stress fluids under thermal convection [70]. Also, recent numerical studies based on lattice Boltzmann models (LBMs) underscored the importance of finite-sized droplets [71, 72, 73], showing that their presence is crucial in enhancing heat flux fluctuations as the volume fraction of the dispersed phase increases, especially by approaching from above the transition from conduction to convection. In addition, unlike Newtonian convection, thermal plumes in non-Newtonian systems can stop before the emergence of other plumes [70]. This arrest can take place after long periods of chaotic oscillations [61], and it may be intimately connected with the finite size of the droplets, as recently suggested by experiments [74] and numerical studies by the authors [73]. Indeed, in Ref. [73], we considered a model yield-stress emulsion and pinpointed a transient intermittent regime in the heat flux. In the latter regime, prolonged conductive periods alternate with quick and intense convective heat bursts triggered by microscopic plasticity and the large spatial correlation of the emulsion; during the heat bursts, coalescence events occur and make the emulsion prone to a phase inversion, whereby the majority phase becomes continuous and the minority phase forms droplets (e.g., an oil-in-water emulsion becomes a water-in-oil emulsion). Such sudden morphological change comes together with a dramatic reduction of the total interface area, which, in turn, determines a drastic variation of the rheological properties [75, 76, 77, 78] and a consequent transition to thermal convection of the phase-inverted emulsion. In general, a phase inversion can be achieved following various cues, such as chemical treatments [79, 80], or mechanical forcing (stirring speed) [81, 78], or via the presence of homogeneous isotropic turbulent conditions [82, 26, 83]. To the best of the authors’ knowledge, except for the study presented in Ref. [73], phase inversion of concentrated emulsions in thermally buoyant flows has not been reported elsewhere.
@@
The scenario portrayed above is rather intriguing and rich. Still, a deep understanding of the behavior of emulsions under thermally driven flows remains elusive. In this work, we significantly advance our understanding of thermally driven emulsions using LBM numerical simulations in the paradigmatic setup of Rayleigh-Bénard thermal convection, i.e., with the emulsion placed between two parallel walls at different temperatures under the effects of buoyancy forces, the latter encoded in the dimensionless Rayleigh number $\operatorname{Ra}$. We systematically change both $\phi$ and $\operatorname{Ra}$ and offer an unprecedented characterization of the emulsion dynamics and a detailed view of the structural changes that the emulsion can experience. As $\operatorname{Ra}$ increases for a given $\phi$, the emulsion enters convective states where structural changes – such as droplet breakup, coalescence, or phase inversion – significantly alter its rheology. Furthermore, we characterize the scaling properties of the droplet size distribution at sufficiently high values of $\operatorname{Ra}$, where two regimes emerge: for low-to-moderate values of $\phi$, the dynamics is dominated by droplet breakup, whereas at higher values of $\phi$, phase inversion becomes predominant.
@@
The paper is organized as follows: the LBM for emulsion modeling is reviewed in Sec. II; dynamical regimes experienced by the emulsions are described in Sec. III; statistical analysis of droplet size distribution at high values of $\operatorname{Ra}$ is provided in Sec. IV. Conclusions will be drawn in Sec. V.
II Numerical modeling of emulsions
@@
We perform numerical simulations of emulsions by employing the open source TLBfind code [84] based on LBMs [85, 86, 87]. To the best of our knowledge, the LBM approach is so far the only one allowing the simulation of realistic emulsion systems with finite-size droplets and non-trivial rheology [88, 89, 90, 91, 26, 71, 72, 92, 73]. The essential features of the numerical methodology implemented in the TLBfind code are here illustrated; more details on the numerical model can be found in [84].
@@
We consider a thermal non-ideal fluid with two components in the framework of the Shan-Chen interaction included in a LBM [93, 94, 33, 95]. The two fluid components have equal viscosity and equal thermal diffusivity. Without loss of generality, hereafter, we refer to “oil (O) phase”, “water (W) phase”, “oil-in-water (O/W) emulsion” and “water-in-oil (W/O) emulsion”. The emulsion dynamics is numerically studied on a two-dimensional lattice by integrating the lattice Boltzmann equation, describing the evolution of the probability distribution function $f_{\beta,i}({\bm{x}},t)$ of finding a fluid particle of the $\beta=\mathrm{O,W}$ phase with kinetic velocity ${\bm{c}}_{i}$ at the discrete time $t$ in the lattice position ${\bm{x}}=(x,z)$. We assume unitary lattice spacing $\Delta x$ and time-lapse $\Delta t$. Kinetic velocities are discretized, and the index $i$ runs over a discrete set of 9 lattice velocities ${\bm{c}}_{i}$ ($i=0,1,\dots,8)$ with constant Cartesian components: ${\bm{c}}_{0}=(0,0)$, ${\bm{c}}_{1,3}=(\pm 1,0)$, ${\bm{c}}_{2,4}=(0,\pm 1)$, ${\bm{c}}_{5,8}=(+1,\pm 1)$, ${\bm{c}}_{6,7}=(-1,\pm 1)$ (D2Q9 LBM scheme). We consider the following lattice Boltzmann equation:
$\small f_{\beta,i}({\bm{x}}+{\bm{c}}_{i},t+1)-f_{\beta,i}({\bm{x}},t)=\Omega_{\beta,i}({\bm{x}},t)+S_{\beta,i}({\bm{x}},t)\ ,$ (1)
@@
where $\tau$ is a characteristic time of the relaxation process. Furthermore, the action of internal and external forces enters via a source term $S_{\beta,i}({\bm{x}},t)$ whose implementation details can be found in [33, 84]. The local equilibrium distribution function $f_{\beta,i}^{(\mathrm{eq})}$ depends on $({\bm{x}},t)$ via the coarse-grained density $\rho_{\beta}=\rho_{\beta}({\bm{x}},t)$ and velocity ${\bf u}={\bf u}({\bm{x}},t)$ fields as
$\small f_{\beta,i}^{(\mathrm{eq})}(\rho_{\beta},{\bf u})=w_{i}\rho_{\beta}\left[1+\frac{{\bf u}\cdot{\bm{c}}_{i}}{c_{s}^{2}}-\frac{{\bf u}\cdot{\bf u}}{2c_{s}^{2}}+\frac{({\bf u}\cdot{\bm{c}_{i}})^{2}}{2c_{s}^{4}}\right].$ (3)
@@
where $\rho=\sum_{\beta}\rho_{\beta}$ is the total density. Eqs. (1) for both components are coupled by reciprocal interaction forces ${\bm{F}}_{\beta}={\bm{F}}_{\beta}({\bm{x}},t)$, embedded in the source term $S_{\beta}$, which encode both inter-component and intra-component interactions, together with external forces. Intra-component forces introduce phase segregation and promote the emergence of interfaces with non-negligible surface tension $\Sigma$ [93, 33]:
@@
where $G_{\beta\beta}^{a}<0$, $G_{\beta\beta}^{r}>0$ are coupling constants and the second summation is performed with dedicated weights $p_{i}$ [33] on a set of links larger than the standard D2Q9 scheme, thus including also next-to-nearest neighbor lattice directions [33, 84]. Inter-component interactions play a fundamental role since they support the simulation of stable emulsions, which differ from simple mixtures of two immiscible fluids. At the hydrodynamic level, LBMs reproduce the Navier-Stokes equations for the hydrodynamic velocity ${\bm{u}}={\bf u}+{\bm{F}}/2\rho$ with kinematic viscosity, $\nu$, related to the relaxation time $\tau$ as $\nu=c_{s}^{2}(\tau-1/2)$ [86, 87]. The Navier-Stokes equations are coupled with the advection-diffusion dynamics of a temperature field $T=T({\bm{x}},t)$. Specifically, in the momentum equation, a buoyancy term in the Boussinesq approximation [97] is added as
${\bm{F}}_{\beta}^{\mathrm{ext}}({\bm{x}},t)=-\rho_{\beta}({\bm{x}},t)\alpha{\bm{g}}T({\bm{x}},t)\ ,$ (7)
where $\alpha$ is the thermal expansion coefficient and ${\bm{g}}=-g\hat{z}$ the gravity acceleration. To reproduce the advection-diffusion dynamics for the temperature field, we introduce an additional probability distribution function, $h_{i}({\bm{x}},t)$, which evolves following a dedicated lattice Boltzmann equation [86, 87]:

### v2

@@ Abstract
Emulsions are paramount in various interdisciplinary topical areas, yet a satisfactory understanding of their behavior in buoyancy-driven thermal flows has not been established. In the present work, we unravel the dynamical regimes of thermal convection in emulsions by leveraging a large set of mesoscale numerical simulations. Emulsions are prepared with a given volume fraction of the initially dispersed phase, $\phi$, ranging from dilute (low values of $\phi$) to jammed emulsions (high values of $\phi$), resulting in different rheological responses of the emulsion, i.e., from Newtonian to non-Newtonian yield-stress behaviors, respectively. We then characterize the dynamics of the emulsions in the paradigmatic setup of the Rayleigh-Bénard convection, i.e., when confined between two parallel walls at different temperatures under the effect of buoyancy forces, the latter encoded in the dimensionless Rayleigh number $\operatorname{Ra}$. We thoroughly investigated the dynamics of the emulsion in the changing of $\phi$ and $\operatorname{Ra}$. For a given $\phi$, at increasing $\operatorname{Ra}$, we observe that the emulsion exhibits convection states, where structural changes may appear (i.e., droplet breakup, coalescence or phase inversion), which inevitably impact the emulsion rheology. For sufficiently high values of $\operatorname{Ra}$, two states of convection are observed: for low/moderate values of $\phi$ (Newtonian emulsions), we observe breakup-dominated dynamics, whereas, for high values of $\phi$ (non-Newtonian emulsions), we observe phase-inverted states. For both scenarios, the droplet size distribution depends on $\operatorname{Ra}$, and scaling laws for the average droplet size are analyzed and quantified. Our results offer new insights into the rich dynamics of emulsions under thermal convection, offering the first detailed characterization of the various dynamical regimes to be expected and their relation with structural changes occurring in such complex fluids.
@@ I Introduction
Emulsions are heterogeneous systems formed by a dispersion of droplets of a liquid phase (e.g., oil) in another immiscible liquid (e.g., water). Adding an emulsifier (e.g., surfactants) stabilizes the dispersion by hindering the coalescence of droplets [1, 2, 3, 4]. Owing to the presence of interfaces that can store part of the injected energy via elastic deformations [5, 6, 7, 8, 9], an emulsion exhibits a variety of rheological responses when stimulated with an external force. An important parameter – discriminating different rheological responses of the emulsion – is the volume fraction, $\phi$, i.e., the ratio between the volume of the dispersed phase and the total volume, which is related to the total interface area for a given number of droplets. For low values of $\phi$, the emulsion behaves as a Newtonian system presenting an augmented effective viscosity that depends on $\phi$ and the viscosity ratio of the two phases [10, 11, 12]. However, for higher values of $\phi$, the rheology of the emulsion is non-Newtonian in that it exhibits a shear-thinning rheology [13], which changes to yield stress behavior when $\phi$ is high enough [1, 14, 6, 7]. If the emulsion is solicited with a sufficiently large external force, droplets can break up and coalesce [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], thus inducing fluctuations of the total interface area and, hence, modifying the rheological response of the system. The application of flowing emulsions in different manufacturing contexts, such as pharmaceutical [28], food [29] or energy industries [30], made these materials the subject of intense scrutiny for decades since a precise control of their stability and rheological properties under specific external forcing and boundary conditions is required in such contexts [31, 32]. Flowing emulsions exhibit highly complex features. Hence, numerical simulations are desirable tools of investigation to address regimes and processes particularly difficult to reproduce in experiments, thus unraveling multiple relevant questions. This aspect is witnessed in the extensive literature on computational studies on droplet/emulsion dynamics, with various methods developed [33, 34, 30, 35, 36, 37]. However, if, on the one hand, considerable effort has been made to investigate the behavior of emulsions under shear, pressure-driven, or even more complex flows (e.g., turbulence), on the other hand, a satisfactory understanding of the behavior of emulsions in buoyancy-driven thermal flows remains incomplete. Thermally driven emulsions have a significant impact in industrial applications, particularly in petrochemical engineering [38], where they influence oil recovery, refining, and transportation processes; but also in food science, where emulsion-based heat transfer is crucial for cooking and processing techniques [39], affecting texture, stability, and nutrient retention. More generally, emulsions may be regarded as “soft materials”, whose behavior in thermal flows is particularly important: in geophysics, for example, these materials are essential for modeling Earth’s mantle convection [40, 41, 42, 43, 44] – where buoyancy-driven currents transport heat from the planet’s interior to its surface, shaping plate tectonics and volcanic activity – and for understanding both explosive and nonexplosive lava flows [45, 46] – where the interaction between molten rock, gas bubbles, and crystals determines magma rheology. Effective analog materials for geodynamic modeling are worth being mentioned [47, 48, 49].
@@
Some computational studies considered droplets (or also bubbles) in convective flows, but they pertain to the case of convective multiphase flows with point-like dispersed objects [50, 51], or multiphase flows where finite-sized droplets/bubbles undergo a dynamics without any stabilization mechanism against their coalescence [52, 53, 54, 24, 55, 56, 57, 58], thus resulting in a system with overall Newtonian rheology. On the other hand, some theoretical/numerical works investigated the role played by the non-Newtonian rheology – and in particular the yield-stress rheology – on the onset of thermal convection using effective single fluid models  [59, 60, 61, 62, 63, 64, 65, 66, 67]. These models rely on the use of the equations of continuum mechanics, supplemented with some non-trivial constitutive relations between the stress and the shear rate to account for non-Newtonian rheology. From the analysis of these effective single fluid models with yield-stress rheology, it has been found that conductive states become linearly stable in the presence of a finite yield stress [59, 60]. Moreover, the system can enter a convective state only after applying a perturbation above a given intensity threshold. The latter increases upon approaching the critical Rayleigh number, marking the transition from conduction to convection [59]. Effective single fluid models, however, neglect the size of microscopic constituents (droplets/bubbles); hence, they cannot account for plasticity and structural changes at mesoscales [68, 69]; the latter, in turn, may have a non-trivial correlation with thermal plumes, as evidenced by some experiments on yield-stress fluids under thermal convection [70]. Also, recent numerical studies based on lattice Boltzmann models (LBMs) underscored the importance of finite-sized droplets [71, 72, 73], showing that their presence is crucial in enhancing heat flux fluctuations as the volume fraction of the dispersed phase increases, especially by approaching from above the transition from conduction to convection. In addition, unlike Newtonian convection, thermal plumes in non-Newtonian systems can stop before the emergence of other plumes [70]. This arrest can take place after long periods of chaotic oscillations [61], and it may be intimately connected with the finite size of the droplets, as recently suggested by experiments [74] and numerical studies by the authors [73]. Indeed, in Ref. [73], we considered a model yield-stress emulsion and pinpointed a transient intermittent regime in the heat flux. In the latter regime, prolonged conductive periods alternate with quick and intense convective heat bursts triggered by microscopic plasticity and the large spatial correlation of the emulsion; during the heat bursts, coalescence events occur and make the emulsion prone to a phase inversion, whereby the majority phase becomes continuous and the minority phase forms droplets (e.g., an oil-in-water emulsion becomes a water-in-oil emulsion). Such sudden morphological change comes together with a dramatic reduction of the total interface area, which, in turn, determines a drastic variation of the rheological properties [75, 76, 77, 78] and a consequent transition to thermal convection of the phase-inverted emulsion. In general, a phase inversion can be achieved following various cues, such as chemical treatments [79, 80], or mechanical forcing (stirring speed) [81, 78], or via the presence of homogeneous isotropic turbulent conditions [82, 26, 83]. To the best of the authors’ knowledge, except for the study presented in Ref. [73], phase inversion of concentrated emulsions in thermally buoyant flows has not been reported elsewhere.
@@
The scenario portrayed above is rather intriguing and rich. Still, a deep understanding of the behavior of emulsions under thermally driven flows remains elusive. In this work, we significantly advance our understanding of thermally driven emulsions using LBM numerical simulations in the paradigmatic setup of Rayleigh-Bénard thermal convection, i.e., with the emulsion placed between two parallel walls at different temperatures under the effects of buoyancy forces, the latter encoded in the dimensionless Rayleigh number $\operatorname{Ra}$. We systematically change both $\phi$ and $\operatorname{Ra}$ and offer a characterization of the emulsion dynamics and a detailed view of the structural changes that the emulsion can experience. As $\operatorname{Ra}$ increases for a given $\phi$, the emulsion enters convective states where structural changes – such as droplet breakup, coalescence, or phase inversion – significantly alter its rheology. Furthermore, we characterize the scaling properties of the droplet size distribution at sufficiently high values of $\operatorname{Ra}$, where two regimes emerge: for low-to-moderate values of $\phi$, the dynamics is dominated by droplet breakup, whereas at higher values of $\phi$, phase inversion becomes predominant.
@@
The paper is organized as follows: in Sec. II, we describe the system set-up and provide a review of the LBM for emulsion modeling; in Sec. III, we analyze the dynamical regimes experienced by the emulsions; in Sec. IV, we provide statistical analysis of droplet size distribution at high values of $\operatorname{Ra}$. Conclusions will be drawn in Sec. V.
II System set-up and numerical modelling
In this section, we provide details on the system set-up and the numerical modeling of emulsions. In Sec. II.1, we describe the Rayleigh-Bénard configuration we used, the protocol for the emulsion preparation, and the relevant observables. In Sec. II.2, a brief review of the employed LBM for the simulations of emulsions is provided.
II.1 System set-up
We consider two fluid components with equal kinematic viscosity $\nu$ and equal thermal diffusivity $\kappa$. Without loss of generality, hereafter, we refer to “oil (O) phase”, “water (W) phase”, “oil-in-water (O/W) emulsion” and “water-in-oil (W/O) emulsion”. We consider O/W emulsions in a 2D Rayleigh-Bénard setup [84, 85, 86]. Using a 2D set-up is mainly driven by computational reasons, in that we need to properly resolve the emulsion droplets and achieve sufficient statistics in a reasonable amount of time. 2D convection is different from 3D convection, but it still captures many of its relevant features [87]. We use a domain of size $L\times H$, with $L\approx 2H$ and wall-to-wall distance $H\approx 20d$, where $d$ is the initial average droplet diameter, which is fixed for all simulations. Two no-slip walls are set at $z=\pm H/2$, and periodic boundary conditions are applied along the $x$ direction. The walls are kept at constant temperatures, $T_{\mathrm{hot}}$ (bottom wall) and $T_{\mathrm{cold}}$ (top wall) (see Fig. 1), corresponding to a temperature jump $\Delta T=T_{\mathrm{hot}}-T_{\mathrm{cold}}$. Gravity acceleration ${\bm{g}}$ is acting in the negative $z$ direction, i.e., ${\bm{g}}=-g\hat{\bm{z}}$. The system is analyzed in terms of continuous fields depending on position ${\bm{x}}=(x,z)$ and time $t$: the hydrodynamical velocity ${\bm{u}}={\bm{u}}({\bm{x}},t)$, the density of the oil (water) phases, $\rho_{\text{O}}({\bm{x}},t)$ ($\rho_{\text{W}}({\bm{x}},t)$), and the temperature field $T({\bm{x}},t)$. In the bulk phases, the dynamics for ${\bm{u}}$ is set by the Navier-Stokes equation with viscosity $\nu$; furthermore, non-ideal interfaces with a non-zero surface tension $\Sigma$ separate regions with majority of water phase ($\rho_{\text{O}}\ll\rho_{\text{W}}$) from regions with majority of the oil phase ($\rho_{\text{W}}\ll\rho_{\text{O}}$). The dynamics for the temperature field $T({\bm{x}},t)$ is governed by an advection-diffusion equation, with the advection set by the hydrodynamical velocity ${\bm{u}}$ and the diffusion regulated by the thermal diffusivity $\kappa$. Further details on the numerical modeling are provided in Sec II.2.
To create O/W emulsions, we follow a dedicated preparation protocol: the oil and water densities are initialized in such a way that an initial number of circular oil droplets, $\mathrm{N}^{\mathrm{init}}_{\mathrm{O}}$, corresponding to the desired volume fraction of the dispersed phase, $\phi$, are arranged in a honeycomb-like configuration. The number of initial droplets ranges from $\mathrm{N}^{\mathrm{init}}_{\mathrm{O}}=220$ for the most dilute emulsion ($\phi=0.16$) to $\mathrm{N}^{\mathrm{init}}_{\mathrm{O}}=700$ for the most packed emulsion ($\phi=0.84$). To create a slightly polydisperse emulsion, we take the honeycomb-like configuration, we add a small random perturbation to the initial centers-of-mass positions of the droplets and the density field of the continuous phase, and then we leave the emulsion free to relax, with no applied buoyancy force, towards its more energetically favorable configuration [88] (see Fig. 1). We compute the volume fraction of the initially dispersed phase, $\phi$, as the fraction of the domain size initially occupied by the oil phase
$\phi=\frac{A_{\mathrm{O}}}{A_{\mathrm{tot}}},$ (1)
where $A_{\mathrm{O}}$ is the area occupied by the oil phase, computed as in Refs. [71, 72], and $A_{\mathrm{tot}}=L\times H$. After this preparation step, buoyancy forces are added. In a homogeneous fluid, an important parameter that controls the various dynamical regimes of thermal convection is the Rayleigh number [89, 90, 85, 91, 86]
$\operatorname{Ra}=\frac{\alpha g\Delta TH^{3}}{\nu\kappa},$ (2)
where $\alpha$ is the thermal expansion coefficient. Other parameter are the Prandtl number $\Pr=\nu/\kappa$, and the aspect-ratio $\Gamma=L/H$. In particular, the conduction-to-convection transition occurs as $\operatorname{Ra}$ exceeds a critical threshold [89]. In the case of emulsions, however, as mentioned earlier, interfaces affect the viscous/elastic-visco-plastic characteristics of the system response to the forcing. Therefore, $\operatorname{Ra}$, as written in Eq. (2), does not carry a uniquely determined dynamical information as $\phi$ increases since the effective viscosity of the emulsion acquires a dependency on the local shear rate and droplet plasticity at mesoscales is enhanced. In other words, since the emulsion is a structured fluid made of droplets at mesoscales, it is not correct to think that the thermal behavior of the emulsion can be obtained from that of a homogeneous system with a different viscosity. For these reasons, in the following, we consider the kinematic viscosity of the corresponding homogeneous system in the definition of both $\operatorname{Ra}$ and $\Pr$, and we keep $\operatorname{Ra}$ as a dimensionless measure of the imposed forcing. Another relevant parameter is the Weber number, defined as $\operatorname{We}=\rho U^{2}\ell/\Sigma$ (where $\rho$ is the emulsion density, $U$ is a characteristic flow velocity and $\ell$ a typical droplet-scale length, e.g., the droplet mean radius). However, since we keep the surface tension $\Sigma$ constant, the $\operatorname{We}$ values are related to the $\operatorname{Ra}$ values (we will provide more details about it in the next section). Similarly to previous studies [24, 55, 56, 57, 58], we do not consider here the temperature dependence of the surface tension. This is indeed justified as in actual expertimental conditions the Marangoni number can be kept small, and then thermocapillary effects are negligible. The modelling of a temperature dependent surface tension and the study of phenomena where this aspect is relevant, are indeed an interesting matter for future works.
Taking all together, the choice of having unitary viscosity and thermal diffusivity ratios, together with imposing no-slip boundary conditions, reduces to four the number of parameters that define the behavior of the emulsion: $\operatorname{Ra}$, $\Pr$, $\Gamma$, and $\phi$. To simplify the problem, we fix $\Pr=1$ and $\Gamma=2$ in all the simulations and explore the various emerging dynamical regimes in the parameter space spanned by $\operatorname{Ra}$ and $\phi$.
where $u_{z}$ is the vertical component of the hydrodynamical velocity and $\langle\dots\rangle_{x,z}$ denotes the space average. Notice that when emulsions reach their statistically steady states, the Nusselt number in Eq. (3) exhibits fluctuations in time around a mean value [71, 72], hereafter indicated as $\overline{\operatorname{Nu}}$ (see also Sec. III.1):
$\overline{\operatorname{Nu}}=\langle\operatorname{Nu}\rangle_{t}$
where $\langle\dots\rangle_{t}$ denotes the time average.
II.2 Lattice Boltzmann modelling of emulsions
@@
Simulations of emulsions are performed by employing the open source TLBfind code [88] based on LBMs [94, 95, 96]. To the best of our knowledge, the LBM approach is so far the only one allowing the simulation of realistic emulsion systems with finite-size droplets and non-trivial rheology [97, 98, 99, 100, 26, 71, 72, 101, 73, 102]. The essential features of the numerical methodology implemented in the TLBfind code are here illustrated; more details on the numerical model can be found in [88].
@@
We consider a thermal non-ideal fluid with two components in the framework of the Shan-Chen interaction included in a LBM [103, 104, 33, 105]. The emulsion dynamics is numerically studied on a two-dimensional lattice by integrating the lattice Boltzmann equation, describing the evolution of the probability distribution function $f_{\beta,i}({\bm{x}},t)$ of finding a fluid particle of the $\beta=\mathrm{O,W}$ phase with kinetic velocity ${\bm{c}}_{i}$ at time $t$ in the position ${\bm{x}}=(x,z)$. We assume unitary lattice spacing $\Delta x$ and time-lapse $\Delta t$. Kinetic velocities are discretized, and the index $i$ runs over a discrete set of 9 lattice velocities ${\bm{c}}_{i}$ ($i=0,1,\dots,8)$ with constant Cartesian components (D2Q9 LBM scheme): ${\bm{c}}_{0}=(0,0)$, ${\bm{c}}_{1,3}=(\pm 1,0)$, ${\bm{c}}_{2,4}=(0,\pm 1)$, ${\bm{c}}_{5,8}=(+1,\pm 1)$, ${\bm{c}}_{6,7}=(-1,\pm 1)$. We consider the following lattice Boltzmann equation:
$f_{\beta,i}({\bm{x}}+{\bm{c}}_{i},t+1)-f_{\beta,i}({\bm{x}},t)=\Omega_{\beta,i}({\bm{x}},t),$ (4)
@@
where $\tau$ is a characteristic time of the relaxation process. The local equilibrium distribution function $f_{\beta,i}^{(\mathrm{eq})}$ depends on $({\bm{x}},t)$ via the coarse-grained density $\rho_{\beta}=\rho_{\beta}({\bm{x}},t)$ and equilibrium velocity $\bar{{\bf u}}_{\beta}=\bar{{\bf u}}_{\beta}({\bm{x}},t)$ fields as
$\small f_{\beta,i}^{(\mathrm{eq})}(\rho_{\beta},\bar{{\bf u}}_{\beta})=w_{i}\rho_{\beta}\left[1+\frac{\bar{{\bf u}}_{\beta}\cdot{\bm{c}}_{i}}{c_{s}^{2}}-\frac{\bar{{\bf u}}_{\beta}\cdot\bar{{\bf u}}_{\beta}}{2c_{s}^{2}}+\frac{(\bar{{\bf u}}_{\beta}\cdot{\bm{c}_{i}})^{2}}{2c_{s}^{4}}\right].$ (6)
@@
where $\rho=\sum_{\beta}\rho_{\beta}$ is the total density. Furthermore, the action of internal and external forces enters via the equilibrium velocity
$\bar{{\bf u}}_{\beta}={\bf u}+\frac{\tau{\bm{F}}_{\beta}}{\rho_{\beta}}$
where the force term ${\bm{F}}_{\beta}={\bm{F}}_{\beta}({\bm{x}},t)$ encodes both inter-component and intra-component interactions, together with external forces, ${\bm{F}}_{\beta}={\bm{F}}^{\mathrm{intra}}_{\beta}+{\bm{F}}^{\mathrm{inter}}_{\beta}+{\bm{F}}_{\beta}^{\mathrm{ext}}$. Intra-component forces introduce phase segregation and promote the emergence of interfaces with non-negligible surface tension $\Sigma$ [103, 33]:
@@
where $G_{\beta\beta}^{a}<0$, $G_{\beta\beta}^{r}>0$ are coupling constants and the second summation is performed with dedicated weights $p_{i}$ [33] on a set of links larger than the standard D2Q9 scheme, thus including also next-to-nearest neighbor lattice directions [33, 88]. The pseudo-potential used in Eq.  (II.2) is set to the original form proposed by Shan & Chen [103], $\psi_{\beta}=\rho_{0}\,(1-\exp(-\rho_{\beta}/\rho_{0}))$. Inter-component interactions play a fundamental role since they support the simulation of stable emulsions, which differ from simple mixtures of two immiscible fluids. At the hydrodynamic level, LBMs reproduce the Navier-Stokes equations for the hydrodynamic velocity ${\bm{u}}={\bf u}+{\bm{F}}/2\rho$ with kinematic viscosity, $\nu$, related to the relaxation time $\tau$ as $\nu=c_{s}^{2}(\tau-1/2)$ [95, 96]. Notice that, although spurious currents are present in the employed LBM [107, 108], they do not affect the observed phenomenology.
The Navier-Stokes equations are coupled with the advection-diffusion dynamics of a temperature field $T=T({\bm{x}},t)$. Specifically, in the momentum equation, a buoyancy term in the Boussinesq approximation [109] is added as
${\bm{F}}_{\beta}^{\mathrm{ext}}({\bm{x}},t)=-\rho_{\beta}({\bm{x}},t)\alpha{\bm{g}}T({\bm{x}},t).$ (10)
To reproduce the advection-diffusion dynamics for the temperature field, we introduce an additional probability distribution function, $h_{i}({\bm{x}},t)$, which evolves following a dedicated lattice Boltzmann equation [95, 96]: